Translate the problem into a pair of linear equations in two variables. Solve the equations using either elimination or substitution. State your answer for the specified variable.
A merchant has coffee worth $3 a pound that she wishes to mix with 6 pounds of coffee worth $8 a pound to get a mixture that can be sold for $4 a pound. How many pounds of the $3 coffee should be used?

Answer by ankor@dixie-net.com(22740)

A merchant has coffee worth $3 a pound that she wishes to mix with 6 pounds of
coffee worth $8 a pound to get a mixture that can be sold for $4 a pound.
How many pounds of the $3 coffee should be used?
:
It does not make sense to use two variables in this problem
:
Let x = amt of $3 coffee:
:
A mixture equation
3x + 8(6) = 4(x+6)
:
3x + 48 = 4x + 24
:
48 - 24 = 4x - 3x
24 = x
:
24 lb of $3 coffee required to make 30 lb of $4 coffee